I am having trouble with inverse operations HELP! is there a trick to memorize the setps?!?! Also
what is the anwer to the foloowing?
12h=60
and
k- 14=29
n divided by 6 = -9???????????????????????????????????????

Answer by longjonsilver(2297)

maths is a language. You have to learn how the language is made up. The actual rules are very few and quite simple and ALWAYS apply.

12h=60 says 12*h=60. This means 12 multiplied by a number (the h) equals 60. Can you figure that answer out just by looking? The answer is 5... 12*5 is 60.

Anyway, mathematically, we have 12*h = 60. We want to know what 1*h is equal to... The question is, we have "12times". How to convert that to "1times"? Answer is "DO THE OPPOSITE OF WHAT YOU HAVE". Opposite operation to multiply is? Divide..so divide by 12.

My Second Rule of algebra... "WHAT YOU DO TO ONE SIDE, YOU MUST DO TO THE OTHER SIDE" since we have an equals sign...something is equal to something else. If you are going to alter one side, then you have to alter the other side too... So, writing every step down, we have:

12*h = 60
(12*h)/12 = 60/12
1*h = 5 (since 12/12 is 1)
or, h=5

What we would tend to write is:
12h=60
h=60/12
h=5

That is the more "fluent" version of the working, but hopefully you can see the similarity.

So, k-14=29. We have "k minus 14". We want the k by itself, so we need to get rid of the "minus 14". We do this by applying the opposite, which is +14 and we have to "do it to both sides"...SAME TWO RULES as before!

k-14=29
k-14+14=29+14
k = 43

n/6 = -9
we have "n divided by 6" So we have to do what? And we must do it to both sides, so... you try doing it

jon.